# RODS: Robust Optimization Inspired Diffusion Sampling for Detecting and Reducing Hallucination in Generative Models

### v1

@@ Abstract
Diffusion models have achieved state-of-the-art performance in generative modeling, yet their sampling procedures remain vulnerable to hallucinations—often stemming from inaccuracies in score approximation. In this work, we reinterpret diffusion sampling through the lens of optimization and introduce RODS (Robust Optimization–inspired Diffusion Sampler), a novel method that detects and corrects high-risk sampling steps using geometric cues from the loss landscape. RODS enforces smoother sampling trajectories and adaptively adjusts perturbations, reducing hallucinations without retraining and at minimal additional inference cost. Experiments on AFHQv2, FFHQ, and 11k-hands demonstrate that RODS improves both sampling fidelity and robustness, detecting over 70% of hallucinated samples and correcting more than 25%, all while avoiding the introduction of new artifacts.
@@ 1 Introduction
To understand how diffusion samplers can fail, it is helpful to view sampling as an optimization process. Consider the probability flow ordinary differential equation (PF-ODE) as guiding a climber down a mountain ridge in dense fog. At each time $t$, the learned score function provides a direction of descent; as time progresses ($t\downarrow 0$), the fog lifts and the terrain becomes clearer. This step-by-step refinement closely resembles the continuation method in numerical optimization, where one solves a sequence of gradually harder problems locally to approach a final objective. However, when the score function is imperfect—as is often the practice case—it can point in misleading directions. In benign regions, such errors are tolerable, but in sensitive areas, they may flip the intended descent direction entirely, sending the trajectory off-manifold and leading to hallucinated outputs saharia2022photorealistic ; balaji2022ediff ; aithal2024understanding .
@@
  • We conduct extensive numerical experiments on high-dimensional, real-world image datasets—including AFHQv2, FFHQ, and a Hand dataset—that go beyond synthetic 1D or 2D Gaussian examples. Our results show that RODS successfully detects over 70% of hallucinated samples and corrects more than 25%, all without introducing new artifacts or requiring any modification to the pre-trained diffusion model.
@@ 3 An Optimization View of Diffusion Sampling Process
In diffusion models, the noise level $t$ plays the role of “smoothing strength.” At a high noise $t$, the convolved distribution $p_{t}(x)\;=\;p_{0}(x)\,*\,\mathcal{N}(0,t^{2}I)$ is more diffuse than $p_{0}(x)$, so its negative log-density $f_{t}(x)\;=\;-\log p_{t}(x)$ is relatively smooth and has fewer local minima than $f_{0}(x)=-\log p_{0}(x)$. As $t$ decreases, $f_{t}(x)$ transitions from a highly smoothed version to the original, sharper landscape. This parallels the continuation principle: one “continues” from an easy subproblem at high noise ($t\approx T$) to the more difficult, exact $-\log p_{0}(x)$ at $t=0$. Such procedure equivalence is formalized in Theorem 1, with the proof detailed in the Appendix A.1.
@@
Moreover, $f_{t}(x)$ shifts continuously from a heavily smoothed version at $t=T$ toward $-\log p_{0}(x)$ at $t=0$, aligning with continuation.
@@
Overall, because diffusion lowers the noise level $t$ in discrete steps, we can view the entire sampling process as a continuation method: each subproblem corresponds to minimizing $f_{t_{i}}(x)$, starting from $t_{N}=T$ (high noise, smooth landscape) and ending at $t_{0}=0$ (true but sharp $-\log p_{0}(x)$). Algorithm 1 details this procedure. It is numerically identical to solving the ODE (5) with an Euler integrator but emphasizes the iterative optimization interpretation. Furthermore, the proximal operator on each subproblem (7) matches the Euler updates, as illustrated in Figure 2. More sophisticated ODE solvers (Heun, RK4) correspond to more elaborate ways of approximating $y^{*}$ in each subproblem
4 Robust Optimization inspired Diffusion Sampler (RODS)
Recent studies have shown that diffusion models sometimes produce unrealistic or “hallucinated” outputs narasimhaswamy2024handiffuser ; borji2023qualitative , especially when sampling paths pass through low-density regions where the model’s estimates become unreliable aithal2024understanding . Standard sampling methods closely follow the model’s estimated direction but lack built-in safeguards against sudden instabilities or sharp changes in behavior.
$\min_{x}\;\max_{\hat{f}_{t}\,\in\,\mathcal{F}_{t}}\hat{f}_{t}(x),$
@@ 4 Robust Optimization inspired Diffusion Sampler (RODS)
where $\mathcal{F}_{t}$ represents a set of plausible variations around the estimated function. This setup asks: what is the worst-case outcome in the surrounding region, and how can we choose $x$ to avoid bad surprises? As a result, we anticipate that, the sampling process becomes more stable, particularly in regions where the model’s approximation is noisy or inconsistent.
Specifically, we monitor local curvature changes—sharp increases in curvature often signal transitions into unstable or poorly approximated zones, where hallucinations are more likely to occur.
@@ 4.1 Curvature Change Detection
To decide when RODS should switch from a standard to a robust update, we need a fast signal that a sampling step is entering unstable territory. As mentioned in song2019generative , generalization errors arise during the diffusion reverse process in low-density regions. This results in the learned score function deviating from the sharp, high-gradient structure of the ground truth score $\nabla_{x}\log P_{t}$, and leads to potential hallucination effects in the diffusion model, where generated samples may deviate from realistic outputs. aithal2024understanding proposed detecting such errors by measuring prediction instability along the temporal direction of the denoising trajectory. In contrast, our approach focuses on identifying instability in the spatial domain. To detect landscapes of the vector field $v(x)$, we propose using an index $\mathcal{H}(x)$, inspired by the maximum eigenvalue of the second derivative of the score norm:
@@
where $\rho$ is the detection radius. Intuitively, $\mathcal{H}(x)$ measures the degree of fluctuation in the magnitude of the score along direction $\delta$. For theoretical insights, please refer to Appendix A.2. We consider the current step $x$ to have entered a high-risk region when $\mathcal{H}(x)>\epsilon$, where $\epsilon$ is a predefined detection threshold. In practice, the choice of $\epsilon$ reflects a trade-off: a lower threshold increases sensitivity and enables the model to detect more hallucinated samples, but may also lead to false positives by flagging benign regions as unstable.
@@ 4.3 RODS Algorithm
Therefore, once we detect that the sampling path is entering a potentially unstable region—typically characterized by sharp curvature changes—we apply a more cautious update to help prevent unrealistic or unstable samples. The full procedure is outlined in Algorithm 2. It closely follows the deterministic sampling style of EDM karras2022elucidating , but selectively activates robust steps when the local geometry suggests the model may be unreliable.
In each iteration, the algorithm first estimates the current score direction and checks whether the score landscape is changing too sharply using the curvature index $\mathcal{H}(x_{i})$. If the region is flagged as high-risk, it searches for the most sensitive direction $\delta_{i}$ that could reveal model uncertainty. This is done by maximizing a function $g_{t_{i}}(x_{i},\delta)$, which can be customized to reflect different types of risk, such as SAS or CAS that we’ve mentioned before.
@@ 5 Experiment
Due to substantial differences across prior works in methodology, assumptions, and evaluation protocols, there is currently no standardized benchmark for hallucination detection, generation and correction. We summarize and contrast related methods in Appendix C. In particular, we perform manual hallucination annotation for evaluation (Appendix B.6).
@@
We generate 1,080 samples from the AFHQv2 dataset using random seeds from 0 to 119, with a batch size of 9. With the robust region :=$\mathcal{M}:=\{\|\delta\|=1\}$ and the detection threshold $\epsilon=0.1$, our method detects 7 of 8 labeled hallucinations, improves 7 unlabeled cases, and introduces no degraded outputs—demonstrating its robustness. As shown in Figure 3, the left panel illustrates successful visual corrections by RODS-CAS, while the right panel summarizes detection and correction performance. Notably, most samples labeled as non-hallucinated by human annotators remain visually unchanged after the robust update (Appendix Figure 13 and 14). In some cases, however, even these non-hallucinated samples exhibit subtle improvements—appearing more semantically coherent after refinement (e.g., top-left in Appendix Figure 8).
@@
We then apply the same evaluation procedure to the AFHQ dataset, with :=$\mathcal{M}:=\{\|\delta\|=8\}$ and $\epsilon=0.09$; an ablation study on threshold sensitivity is provided in Appendix B.2. Our method detects 29 of 40 labeled hallucinations and improves 9 of the detected images without introducing any visible worse cases. As shown in Figure 4, the left panel illustrates representative visual corrections, and the right panel reports quantitative results.
Figure 4: FFHQ sampling results. Left: Visual results across EDM-Euler, EDM-Heun, and our RODS-SAS, RODS-CAS. Hallucinations include distorted eye regions and implausible facial geometry. Right: Confusion matrix at $\epsilon=0.09$ and $\max\|\delta\|=1$ over 1,080 samples using the proposed hallucination index $\mathcal{H}(x)$. RODS-CAS detects 72.5% of labeled hallucinations, corrects 9 cases (better), and introduces no degradation (worse).
@@
To evaluate performance in a more challenging setting, we test on the 11k-hands dataset, where hallucinations are more frequent and severe. We generate 900 samples using seeds 0 to 99 with a batch size of 9. To ensure high sensitivity, we adopt a stricter configuration with :=$\mathcal{M}:=\{\|\delta\|=30\}$ and $\epsilon=0.005$. Under this setting, our method successfully detects 165 out of 169 human-labeled hallucinations. As a trade-off for this high sensitivity, the method attempts to revise every suspicious sample.
Based on ablation studies (Appendix B.3), we observe that critical steps—those most responsible for hallucinations—tend to occur in the middle of the sampling trajectory. To reduce computational cost while preserving performance, we employ a critical-step truncation strategy (Appendix B.5). We compare EDM-Euler, EDM-Heun, and our proposed method RODS-CAS. As shown in Figure 5, RODS-CAS achieves the lowest hallucination rate while correcting the highest proportion of faulty samples, and introduces no new hallucinations.
These results—spanning multiple datasets and varying task difficulty—demonstrate that our method generalizes effectively: it identifies potentially hallucinated samples at a controllable sensitivity; after applied robust corrections to the identified hallucination, it achieves over 25% correction for true positive depending on the difficulty of a given task, and preserves output quality even for false positives.
Figure 5: 11k-hands sampling results. Left: Visual results for different samplers. RODS-CAS corrects hallucinations such as extra or missing fingers while preserving anatomical plausibility. Top-Right: Confusion matrix at threshold = 0.005 over 900 samples using the proposed hallucination index $\mathcal{H}(x)$. Bottom-Right: Quantitative metrics: hallucination rate (↓), correction rate (↑), new hallucination rate (↓), and inference time (in seconds).
@@ 6 Conclusion
In this work, we proposed the equivalence between the diffusion sampling process and the numerical optimization continuation method. Build upon such optimization point-of-view, we introduced RODS, that is to detect and correct hallucinations in generative models. By leveraging geometric information from the loss landscape, our method adaptively adjusts the sampling direction in high-risk regions, significantly improving sample quality without additional training or excessive compute. We validated our approach across three datasets—AFHQv2, FFHQ, and 11k-hands. While RODS-CAS shows promising improvements over standard samplers in reducing hallucinations and correcting semantic inconsistencies, its effectiveness has so far been demonstrated only in relatively unconditional settings. Its generalizability to text-to-image or conditional generation tasks should be straightforward but requires a thorough study in future work.

### v2

@@ Abstract
Diffusion models have achieved state-of-the-art performance in generative modeling, yet their sampling procedures remain vulnerable to hallucinations—often stemming from inaccuracies in score approximation. In this work, we reinterpret diffusion sampling through the lens of optimization and introduce RODS (Robust Optimization–inspired Diffusion Sampler), a novel method that detects and corrects high-risk sampling steps using geometric cues from the loss landscape. RODS enforces smoother sampling trajectories and adaptively adjusts perturbations, reducing hallucinations without retraining and at minimal additional inference cost. Experiments on AFHQv2, FFHQ, and 11k-hands demonstrate that RODS maintains comparable image quality and preserves generation diversity. More importantly, it improves both sampling fidelity and robustness, detecting over 70% of hallucinated samples and correcting more than 25%, all while avoiding the introduction of new artifacts. We release our code at https://github.com/Yiqi-Verna-Tian/RODS.
@@ 1 Introduction
To understand how diffusion samplers can fail, it is helpful to view sampling as an optimization process. Consider the probability flow ordinary differential equation (PF-ODE) as guiding a climber down a mountain ridge in dense fog. At each time $t$, the learned score function provides a direction of descent; as time progresses ($t\downarrow 0$), the fog lifts and the terrain becomes clearer. This step-by-step refinement closely resembles the continuation method in numerical optimization, where one solves a sequence of gradually harder problems locally to approach a final objective. However, when the score function is imperfect—as is often the practice case—it can point in misleading directions. In benign regions, such errors are tolerable, but in sensitive areas, they may flip the intended descent direction entirely, sending the trajectory off-manifold and leading to hallucinated outputs saharia2022photorealistic ; balaji2022ediff ; aithal2024understanding ; oorloff2025mitigating .
@@
  • We conduct extensive numerical experiments on high-dimensional, real-world image datasets—including AFHQv2, FFHQ, and a Hand dataset—that go beyond synthetic 1D or 2D Gaussian examples. Our results show that while maintaining comparable image quality and generation diversity, RODS successfully detects over 70% of hallucinated samples and corrects more than 25%, all without introducing new artifacts or requiring any modification to the pre-trained diffusion model.
@@ 3 An Optimization View of Diffusion Sampling Process
In diffusion models, the noise level $t$ plays the role of “smoothing strength.” At a high noise $t$, the convolved distribution $p_{t}(x)\;=\;p_{0}(x)\,*\,\mathcal{N}(0,t^{2}I)$ is more diffuse than $p_{0}(x)$, so its negative log-density $f_{t}(x)\;=\;-\log p_{t}(x)$ is relatively smooth and has fewer local minima than $f_{0}(x)=-\log p_{0}(x)$. As $t$ decreases, $f_{t}(x)$ transitions from a highly smoothed version to the original, sharper landscape. This parallels the continuation principle: one “continues” from an easy subproblem at high noise ($t\approx T$) to the more difficult, exact $\log p_{0}(x)$ at $t=0$. Such procedure equivalence is formalized in Theorem 1, with the proof detailed in the Appendix A.1.
@@
Moreover, $f_{t}(x)$ shifts continuously from a heavily smoothed version at $t=T$ toward $\log p_{0}(x)$ at $t=0$, aligning with continuation.
@@
Overall, because diffusion lowers the noise level $t$ in discrete steps, we can view the entire sampling process as a continuation method: each subproblem corresponds to minimizing $f_{t_{i}}(x)$, starting from $t_{N}=T$ (high noise, smooth landscape) and ending at $t_{0}=0$ (true but sharp $\log p_{0}(x)$). Algorithm 1 details this procedure. It is numerically identical to solving the ODE (5) with an Euler integrator but emphasizes the iterative optimization interpretation. Furthermore, the proximal operator on each subproblem (7) matches the Euler updates, as illustrated in Figure 2. More sophisticated ODE solvers (Heun, RK4) correspond to more elaborate ways of approximating $y^{*}$ in each subproblem.
@@
4 Robust Optimization inspired Diffusion Sampler (RODS)
Recent studies have shown that diffusion models sometimes produce unrealistic or “hallucinated” outputs narasimhaswamy2024handiffuser ; borji2023qualitative , especially when sampling paths pass through low-density regions where the model’s estimates become unreliable aithal2024understanding . Standard sampling methods closely follow the model’s estimated direction but lack built-in safeguards against sudden instabilities or sharp changes in behavior. Therefore, our proposed hallucination detection builds on the insight that hallucinated behavior is closely correlated with model uncertainty, which can be quantitatively inferred from the local curvature change (rapid fluctuations) of the score field $v(x)=\nabla_{x}\log P_{t}(x)$. For example, sharp increases in curvature proxy often signal transitions into unstable or poorly approximated zones, where hallucinations are more likely to occur.
$\min_{x}\;\max_{\hat{f}_{t}\,\in\,\mathcal{F}_{t}}\hat{f}_{t}(x),$
where $\mathcal{F}_{t}$ represents a set of plausible variations around the estimated function. This setup asks: what is the worst-case outcome in the surrounding region, and how can we choose $x$ to avoid bad surprises? Intuitively speaking, by using RO method, we peek at the place where the model is most uncertain, borrow a direction that could most effectively mitigate potential errors, and use it to update the current one. As a result, we anticipate that, the sampling process becomes more stable, particularly in regions where the model’s approximation is noisy or inconsistent.
@@ 4.1 Curvature Change Detection
To decide when RODS should switch from a standard to a robust update, we need a fast signal that a sampling step is entering unstable territory. As mentioned in song2019generative , generalization errors arise during the diffusion reverse process in low-density regions. This results in the learned score function deviating from the sharp, high-gradient structure of the ground truth score, and leads to potential hallucination effects in the diffusion model, where generated samples may deviate from realistic outputs. aithal2024understanding proposed detecting such errors by measuring prediction instability along the temporal direction of the denoising trajectory. In contrast, our approach focuses on identifying instability in the spatial domain. High local curvature in the score field $v(x)$ serves as an indicator of model uncertainty and, consequently, hallucination risk. To detect landscapes of the score field, we propose using an index $\mathcal{H}(x)$, inspired by the maximum eigenvalue of the second derivative of the score norm:
@@
where $\rho$ is the detection radius. Intuitively, $\mathcal{H}(x)$ measures the degree of fluctuation in the magnitude of the score along direction $\delta$. For theoretical insights, please refer to Appendix A.2. We consider the current step $x$ to have entered a high-risk region when $\mathcal{H}(x)>\epsilon$, where $\epsilon$ is a predefined detection threshold. In practice, the choice of $\epsilon$ reflects a trade-off: a lower threshold increases sensitivity and enables the model to detect more hallucinated samples, but may also lead to false positives by flagging benign regions as unstable. When the surrounding region is reliable (e.g., visually simple), a smaller $\rho$ is preferred, reflecting higher trust in local information. Conversely, a larger $\rho$ makes the approach more conservative, expanding the search when the neighborhood may be misleading.
@@ 4.3 RODS Algorithm
Therefore, once we detect that the sampling path is entering a potentially unstable region, which typically characterized by sharp curvature changes, we apply a more cautious update to help prevent unrealistic or unstable samples. The full procedure is outlined in Algorithm 2. It closely follows the deterministic sampling style of EDM karras2022elucidating , but selectively activates robust steps when the local geometry suggests the model may be unreliable.
In each iteration, the algorithm first estimates the current score direction and checks whether the score landscape is changing too sharply using the curvature index $\mathcal{H}(x_{i})$. If the region is flagged as high-risk, it proceeds through three steps: (1) Search – look around the current sample to find the most sensitive direction $\delta_{i}$ that exposes model uncertainty, obtained by maximizing a function $g_{t_{i}}(x_{i},\delta)$, which can be tailored to different risk criteria (e.g., SAS or CAS); (2) Descent – compute the steepest descent direction $d_{i}$ at the identified worst-case point to neutralize local instability; and (3) Update – move the original sample $x_{i}$ along this safe corrective direction.
@@ 5 Experiment
Due to substantial differences across prior works in methodology, assumptions, and evaluation protocols, there is currently no standardized benchmark for hallucination detection, generation and correction. We summarize and contrast related methods in Appendix C. As a plug-and-play module, RODS can reduce to the default EDM-Euler sampler when the detection threshold $\epsilon$ is high. Therefore, EDM-Euler serves as our baseline method. In particular, we perform manual hallucination annotation for evaluation (Appendix B.6).
@@
We generate 1,080 samples from the AFHQv2 dataset using random seeds from 0 to 119, with a batch size of 9. With the robust region $\mathcal{M}:=\{\|\delta\|=1\}$ and the detection threshold $\epsilon=0.1$, our method detects 7 of 8 labeled hallucinations, improves 7 unlabeled cases, and introduces no degraded outputs—demonstrating its robustness. As shown in Figure 3, the left panel illustrates successful visual corrections by RODS-CAS, while the right panel summarizes detection and correction performance. Notably, most samples labeled as non-hallucinated by human annotators remain visually unchanged after the robust update (Appendix Figure 13 and 14). In some cases, however, even these non-hallucinated samples exhibit subtle improvements—appearing more semantically coherent after refinement (e.g., top-left in Appendix Figure 8).
@@
We then apply the same evaluation procedure to the AFHQ dataset, with $\mathcal{M}:=\{\|\delta\|=8\}$ and $\epsilon=0.09$; an ablation study on threshold sensitivity is provided in Appendix B.2. Our method detects 29 of 40 labeled hallucinations and improves 9 of the detected images without introducing any visible worse cases. As shown in Figure 4, the left panel illustrates representative visual corrections, and the right panel reports quantitative results.
Figure 4: FFHQ sampling results. Left: Visual results across EDM-Euler, EDM-Heun, and our RODS-SAS, RODS-CAS. Hallucinations include distorted eye regions and implausible facial geometry. Right: Confusion matrix at $\epsilon=0.09$ and $\max\|\delta\|=8$ over 1,080 samples using the proposed hallucination index $\mathcal{H}(x)$. RODS-CAS detects 72.5% of labeled hallucinations, corrects 9 cases (better), and introduces no degradation (worse).
Figure 5: 11k-hands sampling results. Left: Visual results for different samplers. RODS-CAS corrects hallucinations such as extra or missing fingers while preserving anatomical plausibility. Top-Right: Confusion matrix at $\epsilon=0.014$ and $\max\|\delta\|=30$ over 900 samples using the proposed hallucination index $\mathcal{H}(x)$. Bottom-Right: Quantitative metrics: hallucination rate (↓), correction rate (↑), new hallucination rate (↓), and inference time (in seconds).
@@
To evaluate performance in a more challenging setting, we test on the 11k-hands dataset, where hallucinations are more frequent and severe. We generate 900 samples using seeds 0 to 99 with a batch size of 9. To ensure high sensitivity, we adopt a stricter configuration with $\mathcal{M}:=\{\|\delta\|=30\}$ and $\epsilon=0.014$. Under this setting, our method successfully detects 169 out of 175 human-labeled hallucinations. As a trade-off for this high sensitivity, the method attempts to revise every suspicious sample.
Based on ablation studies, we found that the key issue lies in the directional error, particularly in low-density, high-uncertainty regions (Appendix B.5). Furthermore, we observe that critical steps (Appendix B.3), those most responsible for hallucinations, tend to occur in the middle of the sampling trajectory. To reduce computational cost while preserving performance, we employ a critical-step truncation strategy (Appendix B.5). We compare EDM-Euler, EDM-Heun, and our proposed method RODS-CAS. As shown in Figure 5, RODS-CAS achieves the lowest hallucination rate while correcting the highest proportion of faulty samples, and introduces no new hallucinations.
To provide a broader generation quality assessment, we have further computed standard image-quality and diversity metrics on 1,000 unconditional samples from FFHQ, AFHQv2, and 11k-hands. Table 1 reports Fréchet Inception Distance (FID) heusel2017gans (which quantifies the similarity between the generated and real image distributions in the Inception feature space), Inception Variance miao2024training (which measures variability in the latent representations), and DreamSim Diversity domingo2024adjoint (which captures the variance of features extracted from a batch of generated images). These results show that RODS-CAS preserves diversity while keeping FID comparable. Moreover, in our generation, the main subjects remain sharp and artifact-free.
Table 1: Comparison of FID, Inception-Var, and DreamSim metrics across datasets, each evaluated on 1k generated samples.
Dataset FFHQ AFHQv2 11k-hands
FID $\downarrow$ Incep-Var $\uparrow$ DreamSim $\uparrow$ FID $\downarrow$ Incep-Var $\uparrow$ DreamSim $\uparrow$ FID $\downarrow$ Incep-Var $\uparrow$ DreamSim $\uparrow$
EDM-Euler 20.48 0.0669 0.407 13.36 0.0539 0.486 14.57 0.0327 0.1427
EDM-Heun 19.56 0.0699 0.422 12.66 0.0548 0.488 14.20 0.0332 0.1476
RODS-CAS 22.73 0.0661 0.400 16.08 0.0542 0.481 16.68 0.0318 0.1509
These results, which spanning multiple datasets and varying task difficulty, demonstrate that our method generalizes effectively: it identifies potentially hallucinated samples at a controllable sensitivity; after applied robust corrections to the identified hallucination, it achieves over 25% correction for true positive depending on the difficulty of a given task, and preserves output quality even for false positives.
@@ 6 Conclusion
In this work, we proposed the equivalence between the diffusion sampling process and the numerical optimization continuation method. Build upon such optimization point-of-view, we introduced RODS, that is to detect and correct hallucinations in generative models. By leveraging geometric information from the loss landscape, our method adaptively adjusts the sampling direction in high-risk regions, significantly improving sample quality without additional training or excessive compute. We validated our approach across three datasets—AFHQv2, FFHQ, and 11k-hands. Overall, RODS-CAS delivers equally diverse, comparably high-quality unconditional samples while reducing hallucinations.
While RODS-CAS shows promising improvements over standard samplers in reducing hallucinations and correcting semantic inconsistencies, several aspects remain open for future study. First, our curvature-based detection provides reliable signals of instability, however, more precise localization could further improve correction accuracy. Empirically, we observe that early corrections tend to influence global semantics, while late ones mainly refine local details. This trade-off between early and late interventions offers flexibility, enabling RODS to adapt its correction strength to different scenarios. Second, our RODS algorithm can be directly implemented within the VE-ODE framework, which benefits from the equivalence between VE-based sampling and the continuation method under the usual parameterization. The extension of RODS to other sampling schemes (e.g., VP or latent-space diffusion) is conceptually straightforward but remains to be validated experimentally. Furthermore, its generalizability to text-to-image or conditional generation tasks should be straightforward but requires a thorough study in future work.